Tier1 Operations Report 2018-07-09

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 9th July 2018

Review of Issues during the week 2nd July to the 9th July 2018.
Current operational status and issues
  • DS upgraded the RAL site connection to JANET to 100gb. The link between the new routers and the older border routers were upgraded from 40gb to 80gb to take advantage of the new site connection speed. However, during this process they appear to have broken IPv6. Although this issue was resolved by 0900 Wednesday morning we experienced a period of time where we has no IPv6 connectivity out of RAL – either via the RAL core or the ‘bypass’ route.
Resolved Castor Disk Server Issues
  • None
Ongoing Castor Disk Server Issues
  • gdss747 - Atlas - d1t0. Drive in intervention
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
  • Worker nodes and CONDOR. We are now running new Docker containers running Singularity. Further changes include an update to use UMD4. This new container is currently being tested and is running on arc-ce04.
  • The new Docker container image (stfc/grid-workernode-c6:2018-07-09.2 and stfc/grid-workernode-c7:2018-07-09.2), has been rolled out across the entire batch farm.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
  • No downtime scheduled in the GOCDB for next 2 weeks
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.

Listing by category:

  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • There is currently a scheduled three day period (July 24th - July 26th), where RAL Tier-1 will be undertaking server room circuit breaker testing. Although it is believed that these tests should not effect our services, such is the nature of these tests all Tier-1 production services should be considered "At Risk" during this time. All VOs are asked to give consideration with respect to possible unexpected outage to any high priority/critical jobs they may wish to run during this period.

GGUS Tickets (Snapshot taken during morning of the meeting).

Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
136028 cms in progress urgent 10/07/2018 11/07/2018 CMS_AAA WAN Access Issues reading files at T1_UK_RAL_Disk WLCG
134685 dteam in progress less urgent 23/04/2018 09/07/2018 Middleware please upgrade perfsonar host(s) at RAL-LCG2 to CentOS7 EGI
124876 ops reopened less urgent 07/11/2016 28/06/2018 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
GGUS Tickets Closed Last week
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
136002 cms solved urgent 09/07/2018 09/07/2018 CMS_Facilities T1_UK_RAL SE Xrootd read failure WLCG
135940 cms solved urgent 04/07/2018 06/07/2018 CMS_Data Transfers Transfers failing to RAL_Disk - no data available WLCG
135901 ops verified less urgent 03/07/2018 06/07/2018 Operations [Rod Dashboard] Issue detected : org.bdii.GLUE2-Validate@site-bdii.gridpp.rl.ac.uk EGI
135822 cms solved very urgent 26/06/2018 05/07/2018 CMS_Central Workflows File Read Problems for Production at T1_UK_RAL WLCG
135740 none closed urgent 20/06/2018 04/07/2018 File Access CVMFS issue on RALfor /cvmfs/dune.opensciencegrid.org/ EGI
135711 cms closed urgent 18/06/2018 06/07/2018 CMS_Central Workflows T1_UK_RAL production jobs failing WLCG
135455 cms closed less urgent 31/05/2018 09/07/2018 File Transfer Checksum verification at RAL EGI
135342 ops verified less urgent 27/05/2018 02/07/2018 Operations [Rod Dashboard] Issue detected : egi.eu.lowAvailability-/RAL-LCG2@RAL-LCG2_Availability EGI
135293 ops verified less urgent 23/05/2018 02/07/2018 Operations [Rod Dashboard] Issues detected at RAL-LCG2 EGI

Availability Report

Target Availability for each site is 97.0% Red <90% Orange <97%
Day Atlas Atlas-Echo CMS LHCB Alice OPS Comments
2018-07-03 100 100 100 100 100 100
2018-07-04 100 100 87 100 100 100
2018-07-05 100 100 100 100 100 100
2018-07-06 100 100 99 100 100 100
2018-07-07 100 100 83 100 100 100
2018-07-08 100 100 89 100 100 96.5
2018-07-09 100 100 100 100 100 100
Hammercloud Test Report
Target Availability for each site is 97.0% Red <90% Orange <97%
Day Atlas HC CMS HC Comment
2018/07/03 100 98
2018/07/04 98 100
2018/07/05 98 100
2018/07/06 100 100
2018/07/07 93 99
2018/07/08 94 99
2018/07/09 100 100

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Notes from Meeting.
  • GGUS ticket this morning ref xrootd - failing request were hitting the external gateway - it shouldn't.
  • Lots of transfers with similar failures. All failing requests hitting xrootd servers are from CMs - None from Atlas.
  • Not enough resource to handle requests.
  • 1. If they talk to the internal gateway will be redirected to the external - though this is not confirmed
  • 2. Mapping within the node/docker container is failing. A failing config somewhere, but TB doesn't think it's our fault as Atlas is running ok for the same config.